### v1

# Token-Aware Coding Flow: A Study with Nano Surge in Reasoning Model thanks: * Corresponding authors

Abstract
With the widespread application of large-scale language models (LLMs) in software engineering, the Chain of Thought (CoT) approach has emerged as a crucial tool for driving automated code generation and optimization. However, despite the significant success of CoT methods in generating high-quality code, the issue of token inflation during the reasoning process remains a formidable challenge to model performance and efficiency, particularly when dealing with complex code smells. Code smells not only affect the maintainability and scalability of code but also significantly increase the computational burden during LLM inference, leading to excessive token consumption and, consequently, reduced reasoning efficiency. This paper introduces an innovative Token-Aware Coding Flow method, aimed at addressing the token inflation problem caused by smelly code in the CoT process. Through experimentation, we validate the synergistic effect of code refactoring and prompt engineering strategies, demonstrating that after eliminating code smells, token consumption during model inference is significantly reduced. The experimental results show that refactored code, while maintaining functional consistency, can reduce token consumption by up to 50%. Additionally, by explicitly prompting the type of code smells in the prompt and incorporating strategies such as context awareness and role constraints, we further optimize the reasoning process, achieving a 24.5% to 30% reduction in token consumption. These optimizations not only significantly enhance the model’s reasoning efficiency and improve code generation quality but also provide new insights for addressing performance bottlenecks in complex code generation tasks.
Index Terms:
code smell, CoT, token consumption
I Introduction
Figure 1: Evolution of software development paradigms. (1) Traditional coding paradigm: human programmers manually translate requirements into code. (2) Augmented coding paradigm: programmers are assisted by base language models (e.g., GPT-3.5, GitHub Copilot) to generate code, reducing manual effort. (3) Reasoning-driven paradigm: advanced reasoning models (e.g., GPT-o3, DeepSeek-R1) interpret requirements and autonomously generate code by simulating human-like thinking, leading to deeper understanding and improved code quality.
With the rapid proliferation of large language models (LLMs), such as GPT-o3 and DeepSeek-R1[1], in the field of software development[2], traditional software development paradigms have undergone profound changes(Fig. 1). In the past, software functionality was implemented solely through manual programming; however, it is now gradually shifting toward human-machine collaboration[3], where code generation and optimization are assisted or directly achieved through the ”Chain of Thought” (CoT) approach[4]. This new development paradigm significantly improves both development efficiency and code quality[5]. However, existing research primarily focuses on LLM performance under ideal conditions, such as code that is well-structured and well-designed[6], neglecting the impact of ”code smells” that are commonly present in industrial settings due to legacy issues or insufficient experience.
Code smells[7] are pervasive in industrial software systems and typically manifest as complex, redundant logical structures, ambiguous naming conventions, excessively long functions or parameter lists, and overly coupled class structures. While these issues may not immediately cause software failures, they severely impair code readability and maintainability in the long run, and may even lead to latent defects[8]. Previous studies have shown that human developers face an additional cognitive burden when dealing with these smelly codes, leading to reduced development efficiency[9]. However, it remains unclear how such smelly code affects model inference efficiency and token consumption, especially in the emerging LLM-driven development paradigm.
Preliminary practical observations suggest that large language models exhibit noticeable ”token inflation”[10] and ”smell regurgitation”[11] phenomena when processing code snippets with smells(Fig. 2). Specifically, the model may repeatedly validate and reason about a particular code fragment or logical unit during inference, resulting in an overall redundant reasoning process that significantly increases token consumption[12]. This not only raises the economic cost of using large models [13] but also severely limits their applicability in resource-constrained environments (e.g., mobile platforms and real-time applications). Therefore, a deeper understanding of and strategies to reduce the negative impact of code smells on model inference efficiency is of great practical significance.
Figure 2: An example of redundant reasoning caused by confusion between C (row-major) and Fortran (column-major) memory layout. The user request involves reshaping a 4D CNN mask into a 2D time series mask compatible with RNNs, which requires correct alignment with Fortran-order expectations. However, the reasoning model repeatedly revisits the same subproblem—how elements are ordered in memory and how reshaping should be done—without converging on a solution. This looped verification, highlighted by repeated statements and backward arrows, illustrates the reasoning inefficiency when the model lacks concrete assumptions about memory layout or reshape semantics.
To further understand and attempt to address the above challenges, this paper proposes a method called Token-Aware Coding Flow. We first constructed a dataset containing ten common Java code smells, as well as a clean code dataset, based on a portion of the CodeXGLUE[14] Text-Code dataset, using the tree_sitter syntax tree construction tool[15]. This dataset was used to observe the token consumption differences when the DeepSeek-R1 model is faced with ”clean code” versus ”smelly code.” The results show that ”smelly code” consumes significantly more tokens per unit time than ”clean code,” highlighting the negative impact of code smells on model inference efficiency[16].
Building on this, we designed a series of exploratory experiments to systematically investigate the impact of different optimization strategies on token consumption. This study focuses on the impact of smelly code on inference models and introduces the ”Token-Aware Coding Flow” method. Through a series of carefully designed empirical experiments, we delve into the token consumption characteristics of smelly code during the Chain of Thought process and propose corresponding optimization strategies[17]. Specifically, this study aims to answer the following research questions (RQs):
  • RQ1: What are the specific impacts of smelly code on the reasoning process of large models (DeepSeek-R1) compared to clean code?
  • RQ2: How does code refactoring (eliminating code smells) impact token consumption during model inference?
  • RQ3: Are there significant differences in token consumption caused by different types of code smells?
  • RQ4: Does explicitly indicating the type of code smells in the model prompt help reduce token consumption during inference?
  • RQ5: In addition to code refactoring, what other effective prompt engineering strategies can further reduce token consumption when reasoning with smelly code?
Through experiments, we obtained the following key findings:
  • Finding1: Compared to clean code, the model consumes approximately 32% more tokens per unit time when processing smelly code.
  • Finding2: After eliminating code smells through refactoring, and ensuring functional consistency, token consumption during inference can be reduced by approximately 50%.
  • Finding3: The impact of different types of code smells varies significantly, with design and structural smells imposing a greater burden on token consumption than naming-related smells.
  • Finding4: Explicitly indicating code smell types in the prompt can reduce token consumption by approximately 24.5%.
  • Finding5: Combining context awareness, role constraints, and cost-sensitive prompt strategies can further reduce token consumption by 2% to 25%.
Based on the above research questions and findings, the main contributions of this paper are as follows:
  • Contribution1: We are the first to explore the impact of code smells on token consumption during model inference.
  • Contribution2: We validate the potential of code refactoring in reducing token consumption and provide a reusable method for automating the refactoring of smelly code.
  • Contribution3: We reveal the differential impact of various types of code smells on the model’s inference process.
  • Contribution4: We propose and validate effective prompt engineering strategies for smelly code handling.
II Context
II-A Code Smell
”Code Smell” refers to code characteristics that, while functional, may indicate underlying issues such as poor quality, maintenance difficulties, and scalability problems. These smells, like duplicated code, long methods, and excessive parameters[18], often result from bad programming practices and can negatively affect performance and readability. While they do not always cause immediate errors, they can complicate future maintenance. Developers often address code smells through refactoring, though manual inspection in large projects can be time-consuming and error-prone. Automated methods using machine learning and pre-trained models are increasingly being explored to improve the detection and elimination of code smells.
II-B Reasoning Module
The reasoning module is essential in intelligent code analysis and generation systems. It uses logical reasoning to interpret input (like code or task descriptions) and generate appropriate outputs. In code generation and repair tasks, it helps the model understand code structure, identify issues, and generate conforming code. With advancements in deep learning, reasoning modules have evolved from rule-based systems to those powered by large pre-trained models like GPT-3 and Codex. These models enhance reasoning accuracy and efficiency, making them crucial for tasks like automated code repair and generation(Fig. 3)[19].
Figure 3: Illustration of the LLM-driven code generation workflow in real-world development scenarios. The process begins with diverse input sources, such as natural language (NL), code translation tasks, bug reports, partial implementations, or test cases. The large language model (LLM) engages in multi-turn reasoning by interpreting the input, thinking step-by-step (e.g., via inner monologue), and generating candidate code. If the generated code fails to meet requirements, the model rethinks the problem by refining prompts and iterating on its reasoning. The process concludes with the production of executable code that satisfies the development goal.
III Experimental Insights
III-A Research Questions
To explore the impact of code smells on the reasoning process of large models, we designed a series of experiments aimed at addressing the following research questions:
  • RQ1: What are the specific impacts of smelly code on the reasoning process of large models (DeepSeek-R1) compared to clean code?
  • RQ2: How does code refactoring (eliminating code smells) impact token consumption during model inference?
  • RQ3: Are there significant differences in token consumption caused by different types of code smells?
  • RQ4: Does explicitly indicating the type of code smells in the model prompt help reduce token consumption during inference?
  • RQ5: In addition to code refactoring, what other effective prompt engineering strategies can further reduce token consumption when reasoning with smelly code?
Through these questions, we aim to gain a deeper understanding of the impact of smelly code on large model inference and explore the potential improvements in reasoning efficiency with different optimization strategies[20]. First, RQ1 primarily focuses on whether smelly code causes large models to consume more tokens during inference, which is crucial for evaluating the model’s ability to handle non-standard code. RQ2 investigates whether refactoring smelly code improves the model’s reasoning efficiency, particularly in terms of reducing token consumption while maintaining functional consistency. RQ3 focuses on the impact of different types of code smells on model reasoning efficiency, helping us understand which types of code smells have a more significant effect on token consumption. RQ4 examines whether explicitly pointing out the type of code smells in the prompt provides the model with more precise contextual information, thereby reducing redundant reasoning steps and lowering token consumption. Finally, RQ5 explores whether other effective prompt engineering strategies, such as context awareness, role constraints, and cost-sensitive strategies, can further optimize reasoning efficiency beyond code refactoring. By addressing these questions, we aim to both validate the specific impact of code smells on large model reasoning performance and provide experimental evidence for improving inference efficiency through refactoring and optimized prompt design.
III-B Tasks
In this study, we designed a series of evaluation tasks to systematically explore the impact of code smells on the reasoning process of large models [21] and to test the effectiveness of different optimization strategies. The following are the specific tasks set for each research question.
III-B1 Observational Experiment
Datasets: We used the Java dataset (Text-To-Code) from CodeXGLUE[14] for the experimental data, which includes 300 samples of smelly code and 300 samples of clean code. These datasets were used to analyze the behavior of the DeepSeek-R1 reasoning model when processing these codes, particularly the differences in token consumption and time consumption.
Metrics: We recorded the total number of tokens consumed by the model for each code and the time taken for reasoning. The token consumption per unit time was calculated as the core metric to assess the efficiency of large models when handling different types of code. Additionally, the generated code was evaluated based on functionality correctness, readability, robustness, and other indicators.
III-B2 Refactoring[22] and Optimization Experiment
Datasets: In this experiment, we used the dataset based on smelly code and applied DeepSeek-R1 for code refactoring to generate refactored code (rf_code). The refactored code was used for chain processing to compare its reasoning efficiency with the original smelly code and clean code.
Metrics: We recorded the number of tokens consumed in each processing step and calculated the Halstead complexity and cyclomatic complexity. By assessing the token consumption per unit of Halstead complexity and per line of code, we evaluated the efficiency improvements of the refactored code during model inference. Further, the accuracy of the generated code was assessed through functionality consistency scores (e.g., CodeBLEU) to ensure that the refactored code maintains functional consistency with the original code.
III-B3 Code Smell Type Comparison Experiment
Datasets: We selected code samples from 10 different types of code smells in the smelly code dataset, with 50 samples from each type, totaling 500 samples. These samples were used to analyze the impact of different types of code smells on token consumption during model inference.
Metrics: By comparing the token consumption during reasoning for different types of code smells, we calculated the token consumption per unit time for each smell type. The differences in the impacts of design, structural, and naming-related smells were further analyzed. All generated code was evaluated for functional consistency to ensure that the model-generated code met the expected functionality during reasoning.
III-B4 Explicit Smell Type Prompt Experiment
Datasets: This experiment was conducted on the previously used smelly code dataset, primarily comparing the token consumption differences when explicitly indicating code smell types in the prompt versus not indicating them.
Metrics: We recorded the number of tokens consumed in each experimental step and calculated token consumption per unit time. After generating the code, functionality consistency was assessed using metrics like CodeBLEU. Additionally, the effect of explicitly indicating code smell types was evaluated based on the model’s reasoning time, token consumption, and output quality.
III-B5 Prompt Optimization Strategy Experiment
Datasets: In this experiment, we tested different prompt optimization strategies on the smelly code dataset[23], analyzing the effects of Context Awareness, Responsibility Tuning, and Cost-Sensitive strategies.
Metrics: We recorded token consumption during the reasoning process for each strategy and calculated the Halstead complexity and cyclomatic complexity for each code. By comparing token consumption per unit complexity and per line of code, we assessed the impact of the three prompt optimization strategies on reasoning efficiency. All experiments were also evaluated for functionality consistency and code generation quality to ensure the generated code was accurate[16].
III-C Implementation
This study’s experimental implementation is built upon existing LLMs and open‑source code‑refactoring tools, with all tasks orchestrated via API calls to guarantee reproducibility and efficiency[24]. Specifically, we employ DeepSeek‑R1 as the inference engine and integrate Tree‑Sitter for code‑smell detection and automated refactoring. The implementation proceeds as follows(Fig. 4):
Figure 4: Experiment Design Process. The diagram shows the steps involved in extracting requirements, generating code, refactoring code, and comparing semantic and code similarities. The flow demonstrates the interaction between different components, including code refactoring and requirement extraction, with the evaluation steps represented through similarity comparison.
  • Data Processing [20] and Smell Annotation: Java code was extracted from the Text‑Code subset of CodeXGLUE and parsed with Tree‑Sitter to detect and classify instances of code smell. Leveraging Tree‑Sitter’s high‑performance syntax analysis, we annotated ten common smell categories, thereby constructing parallel “smelly” and “clean” code datasets. For each smelly sample, we invoked the DeepSeek‑R1 API to execute the reasoning task, recording both token consumption and inference latency. We then applied DeepSeek‑R1’s refactoring API to eliminate smells—producing rf_code—and compared the inference efficiency (tokens consumed and time taken) between original and refactored versions.
  • Prompt Engineering and Optimization[25]:To further reduce token overhead, we evaluated three prompt‑level strategies: context awareness, role constraints, and cost sensitivity. Each strategy was realized by modifying the model’s input prompt, thereby influencing the reasoning trajectory and token usage. For every experimental condition, we issued the corresponding API calls to adjust generation parameters and logged token counts, inference durations, and the resulting code quality.
  • Automation, Logging, and Analysis:All API interactions were fully automated via Python scripts, ensuring consistent and efficient execution. Each experiment’s outputs—including raw logs of token usage, timing, and generated code—were aggregated and processed by custom analysis scripts. These scripts perform data cleansing, statistical aggregation, and visualization, enabling systematic comparison across differing experimental configurations.
IV Experimental Results
IV-1 RQ1: Smelly Code vs. Clean Code
Experimental Design.
We randomly sampled 30 instances from each of the ten individual code‑smell datasets and aggregated them into a single file, smelly_code.jsonl. We then randomly selected 300 examples from the clean‑code dataset, producing clean_code.jsonl. Because the two collections differ substantially in content and structure, we employ the Time‑Scaled Token Consumption metric to evaluate token expenditure during inference. Using the DeepSeek‑R1 model, each sample is assessed across six dimensions—functional correctness, readability, robustness, maintainability, extensibility, and security—while we record the inference procedure, evaluation outcomes, and end‑to‑end execution time. From these logs, we compute the total tokens consumed during both the reasoning and output‑generation phases.
Analysis.
In our analysis, we first compare token consumption per unit time between clean code and smelly code. Directly contrasting absolute token counts would misrepresent efficiency, since the code samples vary greatly in complexity and structure. By normalizing to tokens per unit time, we obtain a more precise measure of model throughput across code types. Statistical evaluation reveals that smelly code consistently demands higher token expenditure(Fig. 5)(Tab. I). This elevated consumption stems from the increased complexity of smelly code: the model must undertake additional reasoning steps and verifications to resolve intricate structural and logical relationships. Conversely, clean code—with its clearer organization—allows the model to complete inference with fewer redundant operations.
Figure 5: Comparison of time-scaled token consumption between clean code and smelly code.
TABLE I: Descriptive Statistics of Time-Scaled Token Consumption
Statistic Clean Code Smelly Code
Count 300 300
Mean 24.44 33.20
Std. Dev. 12.35 18.63
Min 0.61 0.83
25% Quartile 18.55 25.30
Median (50%) 28.13 33.30
75% Quartile 31.68 41.90
Max 58.97 120.16
Findings – Smelly Code vs Clean Code The results show that smelly code leads to significantly higher token consumption during inference than clean code. This is mainly attributed to structural complexity such as redundant logic and poor naming, which forces the model to perform repeated validations. These findings suggest that code smells not only harm maintainability but also increase the inference cost of large models, highlighting the importance of code quality for efficient model reasoning.
IV-A RQ2: Impact of Code Refactoring on Token Consumption under Functional Consistency Constraint
Experimental Setup.
To answer RQ2, we designed an experiment to investigate how code refactoring—specifically the elimination of code smells—affects token consumption during model inference. We first used the DeepSeek‑R1 inference model on both smelly_code.jsonl and clean_code.jsonl to establish baseline token‑usage data. Next, we applied automated refactoring to the smelly code samples, removing common smells (e.g., long methods, duplicated logic) to generate a refactored code set (rf_code). We then re‑ran inference on these refactored samples and compared their token consumption to the original smelly code. To ensure functional consistency, we employed CodeBLEU and other metrics to verify that the refactored code preserved the original behavior.
Analysis.
During analysis, we recorded both the inference duration and token consumption for each trial and computed token consumption per unit time. Under the functional consistency constraint, refactored code is expected to be leaner, reducing unnecessary complexity and redundant patterns. Our results show that after refactoring, the DeepSeek‑R1 model performs inference more efficiently: despite maintaining equivalent functionality(Tab. III), the streamlined code requires significantly fewer tokens(Fig. 6)(Tab. II). In particular, refactoring complex smells—such as excessively long functions and duplicated segments—substantially reduced the number of reasoning steps and validations the model must perform.
Figure 6: Comparison of complexity-normalized token consumption before and after code refactoring.
TABLE II: Descriptive Statistics of Token per Complexity
Statistic Original Code Refactored Code
Count 300 300
Mean 0.1015 0.0576
Std. Dev. 0.1732 0.0816
Min 0.0020 0.0020
25% Quartile 0.0144 0.0094
Median (50%) 0.0334 0.0239
75% Quartile 0.1180 0.0723
Max 1.5613 0.5836
TABLE III: Summary of CodeBLEU and Docstring Similarity Statistics (Mean Values Only)
Category Metric Mean
CodeBLEU code & rf code 0.5224
rf code & rf gc code 0.1803
Docstring code & rf code 0.8604
code & rf gc code 0.7356
rf code & rf gc code 0.7364
Findings – Refactoring Impact Code refactoring yields a substantial reduction in token consumption during inference while preserving functional correctness. Specifically, the refactored code consumes approximately 50% fewer tokens compared to the original smelly code. By removing redundant structures and simplifying complex fragments, refactoring not only enhances code quality but also markedly improves inference efficiency. These results suggest that systematic elimination of code smells is a critical strategy for reducing computational overhead in large‑model reasoning.
IV-B RQ3: Impact of Different Types of Code Smells on Token Consumption
Experimental Design / Setup
To answer RQ3, we designed an experiment to investigate how different types of code smells affect token consumption in the DeepSeek‑R1 inference model. We extracted from smelly_code.jsonl ten common smell categories—complicated_regex_expression, parameter_list_too_long, binary_operator_in_name, complicated_boolean_expression, etc.—with 30 code snippets per category (300 total). We ran DeepSeek‑R1 inference separately on each group, recording the tokens consumed for each snippet. By computing the tokens consumed per unit time for each smell type, we identified which categories impose the greatest inference burden.
Analysis
The first graph(Fig. 7) and the table(Tab. IV))present the average growth rate of token consumption across various code smell types. It is evident that most of the code smells, such as binary_operator_in_name, complicated_boolean_expression, and complicated_regex_expression, lead to a significant increase in token consumption. These types of smells are often characterized by complex logical constructs that demand deeper reasoning from the model, resulting in more tokens being processed[26]. Conversely, the smells func_name and too_long_parameter_list have a relatively lower impact on token growth. This suggests that these types of smells are simpler and less likely to introduce complexity in reasoning, causing fewer tokens to be consumed.
The second graph(Fig. 8) categorizes code smells into four groups: Naming, Expression, Structure, and Design. The Expression category stands out with the highest average token consumption growth, followed by Structure. These categories typically involve complex logical constructs or convoluted code structures, which require more tokens for the model to parse and reason about. In contrast, the Naming and Design categories exhibit lower token consumption growth. Issues like poor naming conventions and design flaws might affect code readability and maintainability, but they do not significantly complicate the reasoning process, leading to lower token usage by the model.
Figure 7: Average token growth rate per code smell type.
Figure 8: Average token growth rate by code smell category.
TABLE IV: Average Token Growth Rate by Smell Type
Smell Type Avg. Growth Rate
binary_operator_in_name 0.3424
complicated_boolean_expression 0.3361
complicated_regex_expression 0.3221
cyclomatic_complexity 0.3285
func_name 0.1592
loops 0.3111
mutation_too_much 0.3321
too_long_parameter_list 0.1942
primitive_obsession 0.2963
too_long 0.3089
TABLE V: Comparison of Code Equivalence Before and After Refactoring (Mean Values Only)
Category Metric Mean
Source Code code & gc code 0.1659
Refactored Code code & rf code 0.5208
rf code & rf gc code 0.1773
Source Code Similarity code & gc code 0.7470
Refactored Code Similarity code & rf code 0.8610
code & rf gc code 0.7295
rf code & rf gc code 0.7340
Findings – Smell Type Impact Our study, conducted with a 70% code functionality similarity threshold(Tab. V), reveals significant variance in token consumption across different code smell categories. Complex code smells—namely complicated_regex_expression, parameter_list_too_long, and complicated_boolean_expression—incur the highest token costs due to increased code complexity and the need for additional reasoning steps. Simpler smells, such as binary_operator_in_name and func_name, result in comparatively minor token overhead. These findings suggest that prioritizing the removal of design and structural smells can maximize inference efficiency and substantially reduce computational overhead.
IV-C RQ4: Impact of Explicitly Indicating Code Smell Types in Prompts on Token Consumption[27]
Experimental Design / Setup
To address RQ4, we designed an experiment to investigate whether explicitly annotating code smell types in the model prompt effectively reduces token consumption during inference. We compared two conditions on the same smelly_code.jsonl dataset: (1) No annotation, using the raw code as input; and (2) Explicit annotation, where the prompt was augmented with statements such as “the code contains complex regular expressions” or “the code has an excessively long parameter list.” For each condition, we invoked DeepSeek‑R1 to perform inference and recorded both total token usage and inference time. We then computed token consumption per unit time to facilitate a direct comparison between annotated and unannotated prompts.
Analysis
Our study, conducted with a 70% code functionality similarity threshold (Tab. VII). During analysis, we examined the token consumption profiles under each prompting condition. Results (Fig. 9)(Tab. VI) indicate that adding explicit smell annotations to the prompt yields a more efficient reasoning process, with a marked reduction in token usage. Explicit annotations supply the model with rich contextual cues about potential issues, enabling it to bypass redundant reasoning and validation steps. In contrast, without smell hints, the model must engage in multiple verification cycles to detect and interpret hidden code complexities, thereby inflating token consumption.
Figure 9: Comparison of total token consumption for smelly code with and without explicit prompt tips
TABLE VI: Descriptive Statistics of Total Token Consumption for Smelly Code (With vs. Without Prompt Tips)
Statistic No Tips With Tips
Count 300 300
Mean 5876.49 4431.19
Std. Dev. 3172.77 2578.35
Min 1095.00 981.00
25% Quartile 3419.50 2312.50
Median (50%) 5351.00 3685.00
75% Quartile 7985.75 5820.00
Max 16440.00 15045.00
TABLE VII: Comparison of CodeBLEU and Similarity Metrics with and without Tips
Metric Mean (With Tips) Mean (No Tips)
CodeBLEU 0.1585 0.1665
Code Similarity 0.7248 0.7619
Findings – Explicit Smell Prompt Explicitly indicating code smell types in the model prompt reduces token consumption by approximately 24.5% compared to unannotated prompts. The provision of clear smell-specific cues allows the model to focus its reasoning on relevant code fragments, eliminating unnecessary validation steps and lowering computational overhead. This strategy thus significantly enhances inference efficiency and mitigates resource usage in large‑model code reasoning.
IV-D RQ5: Prompt Engineering Strategies beyond Refactoring
TABLE VIII: Mean Values of Key Metrics under Different Strategies
Strategy Comp‑Norm Code Line‑Scaled Code Comp‑Norm Token Line‑Scaled Token
Base 0.1015 46.7589 0.6026 277.7973
Context 0.0858 39.3372 0.4979 229.0921
Func 0.0891 40.9848 0.5357 241.2648
Total 0.0981 44.9317 0.4760 220.8598
DevOps 0.0852 38.5346 0.4832 218.3642
QAer 0.0868 40.4725 0.5362 238.8191
SEer 0.0771 35.5723 0.5535 241.9891
AbsCost 0.0815 37.1547 0.4746 202.5059
RelCost 0.0859 38.7631 0.4038 178.0056
Comb1 0.1001 44.5918 0.5705 245.4126
Comb2 0.0981 44.1189 0.5301 227.1736
TABLE IX: CodeBLEU and Docstring Similarity Statistics
Category Method Mean CodeBLEU Mean Code-and-GC-Code Similarity
Baseline - 0.1803 0.7364
Context Awareness Code Context 0.1783 0.7428
Code Function 0.2283 0.7668
All Information 0.2464 0.7828
Responsibility Tuning SEer 0.1593 0.7210
QAer 0.1590 0.7369
DevOps 0.1570 0.7275
Cost Sensitive Absolutely 0.1605 0.7011
Relatively 0.1799 0.6988
Combination All Methods 0.2513 0.7635
All Methods 0.2505 0.7665
Experimental Design / Setup
To address RQ5, we designed a series of experiments(Fig. 10) aimed at evaluating three prompt engineering strategies—Context Awareness, Responsibility Tuning, and Cost Sensitivity—and their effect on reducing token consumption during the model’s reasoning over smelly code. Each strategy modifies the input prompt to guide the model’s cognitive process, driving a more efficient and focused inference, while minimizing unnecessary token expenditure. Specifically:
  • Context Awareness: By infusing the prompts with contextual details such as file path, module or function names, and surrounding code snippets, we enhanced the model’s comprehension of the code’s role within the broader system. This enriched context directs the model’s focus to key segments of the code, thereby avoiding redundant inference over irrelevant portions.
  • Responsibility Tuning: Through the strategic assignment of roles—such as “software engineer,” “QA engineer,” or “DevOps engineer”—we framed the model’s reasoning process to align with a specific perspective. This role-induced focus curtailed extraneous reasoning steps and unnecessary pathways, leading to a more efficient processing of the code.
  • Cost Sensitivity: We introduced explicit limits on the length of the generated output or the token count (e.g., maximum token budget). By constraining the token size, we controlled the model’s computational cost, particularly when generating larger code segments, effectively reducing token expenditure without sacrificing the integrity of the generation.
All experiments were performed on the same dataset of 300 smelly code snippets (smelly_code.jsonl). For each condition, we invoked DeepSeek‑R1, recording the total token consumption and calculating token consumption per unit time. Results were compared against a baseline prompt, which did not include any specific optimization strategies.
Analysis
We analyzed token consumption across different prompt conditions by calculating tokens per unit time for each strategy. Our key observations include(Tab. VIII):
  • Context Awareness: By enriching the prompts with contextual information, the model was able to bypass irrelevant logic and focus on the core components of the code. On average, token consumption decreased by 15–20%, with more significant reductions observed in longer, more complex code fragments.
  • Responsibility Tuning: Assigning roles to the model significantly honed its focus on task‑relevant reasoning steps. This strategy yielded a 10–15% reduction in token consumption, particularly in large‑scale codebases, where the model’s focus was sharpened by the specified role.
  • Cost Sensitivity: Imposing limits on output length or token budget led to a 20–30% reduction in token expenditure. However, overly stringent constraints sometimes led to the truncation of essential code elements, resulting in reduced functional correctness and lower accuracy in the output.
Figure 10: Nano Surge on Token Consumption
Findings – Prompt Optimization Strategies Context Awareness and Responsibility Tuning emerged as the most effective strategies for reducing token consumption beyond refactoring(Tab. IX). Context Awareness decreased token usage by 15–20%, especially on lengthy or complex code. Responsibility Tuning achieved a 10–15% reduction by framing the model’s reasoning role. Cost Sensitivity, while reducing tokens by 20–30%, risked incomplete code generation under strict limits. These results suggest that combining context-rich prompts with precise role framing offers substantial efficiency gains without compromising functionality. Furthermore, the optimizations applied through these strategies also led to improvements in code quality metrics. Specifically, the codeBLEU score increased by 38%, and function similarity was enhanced by 4%, compared to the baseline performance. These improvements indicate that not only were token consumption and reasoning efficiency optimized, but the overall quality of the generated code also saw significant gains, further validating the efficacy of these prompt engineering strategies.
V Related Work
Over recent years, with the widespread application of large language models (LLMs) in natural language processing (NLP)[3], automated code generation and optimization have also made significant strides[5]. Research in this area can be broadly categorized into three directions:
V-A Code Smell Detection and Refactoring
Code smells are syntactic or structural code features that do not immediately impair program functionality but degrade readability, maintainability, and extensibility[7][8][9]. Tufano et al. (2017)[28] proposed an automated smell detection approach by analyzing code structures and patterns to identify potential smell instances. Lanza (2003)[29] further investigated the impact of code smells on program maintainability and introduced corresponding refactoring strategies. Many refactoring tools combine rule‑based and machine learning techniques to automatically detect and repair smells at scale[7]. Nevertheless, efficiently handling complex smells remains challenging, particularly in deep learning–based inference where token consumption and computational overhead can escalate.
V-B LLM‑based Code Generation and Optimization
The advent of pre‑trained language models has spurred extensive research into LLM‑driven code generation. Models such as OpenAI’s GPT series and Codex, pre‑trained on large code corpora, demonstrate strong capabilities in code completion, synthesis, and repair. The Transformer architecture of Vaswani et al. (2017) laid the foundation for these tasks, and subsequent variants such as BERT and GPT have further advanced code generation quality. While LLMs capture syntactic and semantic patterns effectively, their inference often suffers from token inflation, especially when processing complex or smelly code. To mitigate this, researchers have explored integrating automated refactoring with LLM training to jointly optimize code quality and inference efficiency.
V-C Integration of Code Quality and Inference Efficiency
Recently, there has been growing interest in coupling code quality improvements with inference efficiency in pre‑trained models. Pre‑trained models like DeepCode and CodeBERT exhibit strong performance in code quality assessment and automated repair. Empirical studies indicate that higher code quality not only enhances maintainability but also reduces inference costs. For example, Zhang et al. (2020) proposed a deep learning–based optimization method that produces functionally equivalent code while significantly lowering token consumption during inference. Building upon this work, our study investigates how code smell removal and prompt engineering jointly reduce token overhead, particularly on complex code fragments.
VI Conclusion and Future Work
This study investigated strategies to optimize the reasoning pipeline of large language models (LLMs) in order to reduce token consumption when processing code with smells[16][14]. We proposed a combined approach of automated code refactoring and prompt engineering[22], and empirically demonstrated that these strategies significantly enhance inference efficiency[13]. Our experimental results show that code smells substantially inflate token usage during inference, whereas eliminating smells via refactoring and refining prompts markedly reduces computational overhead[8][9]. Specifically, code refactoring alone decreased token consumption by approximately 50%, explicit smell annotations in the prompt yielded a 24.5% reduction, and prompt engineering techniques such as Context Awareness and Responsibility Tuning provided further savings.
Despite these advances, several limitations and avenues for future work remain. First, our experiments focused exclusively on Java and the DeepSeek‑R1 model; future research should extend our methodology to other programming languages and inference engines to assess its generalizability. Second, while refactoring and prompt optimizations proved effective, complex code‐smell scenarios may still challenge inference efficiency[13]. Enhancing model reasoning capabilities for diverse programming tasks and more intricate smell types warrants deeper investigation.
Moreover, current work has largely concentrated on token consumption and inference speed[19]. Future studies should explore the trade‐off between generation quality and efficiency—for example, how to maintain code correctness, maintainability, and extensibility while further reducing token usage[27]. Integrating automated refactoring with deep learning–based optimization could also offer practical guidance for improving code quality in large codebases and advancing intelligent programming tools.
In summary, this research provides new insights into token optimization in code generation and reasoning, highlighting the pivotal role of smell elimination and prompt refinement. As intelligent coding assistants evolve, we anticipate that more granular reasoning optimizations and strategic prompt designs will enable next‐generation code generation systems to meet real‐world demands with greater efficiency and precision.